# Spatial correlations in SIS processes on regular random graphs

### v1

@@ Abstract
In network-based SIS models of infectious disease transmission, infection can only occur between directly connected individuals. This constraint naturally gives rise to spatial correlations between the states of neighboring nodes, as the infection status of connected individuals becomes interdependent. Although mean-field approximations are commonly invoked to simplify disease forecasting on networks, they fail to account for these correlations by assuming that infectious individuals are well-mixed within a population, leading to inaccurate predictions of infection numbers over time. As such, the development of mathematical frameworks that account for spatially correlated infections is of great interest, as they offer a compromise between accurate disease forecasting and analytic tractability. Here, we use existing corrections to mean-field theory on the regular lattice to construct a more general framework for equivalent corrections on regular random graph topologies. We derive and simulate a system of ordinary differential equations for the time evolution of the spatial correlation function at various geodesic distances on random networks, and use solutions to this hierarchy of ordinary differential equations to predict the global infection density as a function of time, finding good agreement with corresponding numerical simulations. Our results constitute a substantial development on existing corrections to mean-field theory for infectious individuals in SIS processes and provide an in-depth characterization of how structural randomness in networks affects the dynamical trajectories of infectious diseases on networks.
@@ I Introduction
The Susceptible-Infected-Susceptible (SIS) model, first introduced by Kermack and McKendrick [1], is a foundational framework for describing the spread of infectious diseases in which individuals do not acquire lasting immunity following infection (e.g. the common cold or many sexually transmitted diseases [2, 3, 4, 5]). In the model, members of a population transition repeatedly between two states: susceptible (S) and infected (I). Many generalized SIS models assume that susceptible and infected populations are well-mixed spatially, such that the rates of contact between infected and susceptible individuals are assumed to be uniform across a population. Such models are often referred to as mean-field approximations, and, while they are highly amenable to mathematical analysis [6, 7, 4], they are generally inappropriate for real-world scenarios, in which infected individuals are spatially ordered or clustered. In such contexts, the dynamical trajectory of a disease is highly sensitive to the spatial distribution of infected individuals, where infection is spread more rapidly when infected individuals are distributed more evenly in space.
@@
Figure 1: (a) Schematic of the SIS model on networks. Infected nodes recover with rate $\gamma$ (conventionally, time is rescaled as $t\to\gamma t$, implying $\gamma=1$) and infect their nearest neighbors, each at rate $\beta$. (b)-(e) Snapshots of SIS model simulations on RRGs in the quasi-steady state. For lower values of average node degree $k_{0}$, the clustering of infectious nodes (orange) is more pronounced than on networks with a higher node degree. Moreover, increasing the infection rate, $\beta$, also decreases the clustering of infectious nodes but increases the overall infection density in the steady state. (f) Steady state infection density on the RRG as a function of average node degree, $k_{0}$, for simulations on RRGs with $N=10^{4}$ nodes, $\beta k_{0}=1.7$, and $\gamma=1$. The dashed line is the expected mean-field value of $1-1/(\beta k_{0})\simeq 0.41$. The initial conditions of all simulations are an initial density of $x_{I}=10^{-2}$ with $100$ infective individuals placed randomly on the network, and all results are averaged over 100 simulations.
One solution to mitigate the inaccuracies of mean-field models is to represent the SIS process on a network, in which nodes represent individuals in a population and edges represent connections between individuals (that is, potential routes for the flow of infection). The dynamics of the SIS model on networks are governed by two processes – the infection per connection between infected and susceptible individuals, occurring at a rate $\beta$, and the recovery of infectious individuals at a rate $\gamma$ (see Fig. 1a).
A significant advance in the mathematical modeling of epidemic dynamics on networks was the development of pairwise approximation methods. In contrast to traditional mean-field models, pairwise models explicitly track the densities of connected node pairs of various types, capturing the essential local correlations that arise from network structure and disease transmission mechanisms. Such models established the formulation and analytic power of pairwise ODE frameworks, demonstrating their improved accuracy in regimes where local correlations are significant, such as networks with low degree or heterogeneous topologies [8, 9, 10].
Building on this foundation, modeling SIS processes on a network leads to the emergence of spatial correlations between infectious individuals, since transmission can only occur via direct connections to currently infected neighbors (see Figs. 1b-e). This local transmission mechanism leads to epidemic trajectories that differ markedly from those predicted by mean-field models, especially on networks with lower average node degree, where these deviations are more pronounced (Fig. 1f), or regular lattices. Existing corrections to mean-field theory offer an ordinary differential equation (ODE)-based approach for forecasting disease dynamics, relying on closure approximations that assume spatial correlations between infectious individuals decrease rapidly with shortest path distance [8, 11, 12]. Similarly, for networks with higher degrees of topological randomness—such as random perturbations on the regular lattice achieved through stochastic edge swapping algorithms—mean-field theory is also inaccurate; yet, no analytical framework exists to predict the spread of infection. Consequently, analyzing SIS processes on such networks remains challenging, and simulation-based approaches are typically preferred over analytical methods for predicting disease spread.
@@
To address this problem, this work extends previous corrections to mean-field theory to derive a general framework for the accurate forecasting of disease spread in SIS processes on random network topologies. Motivated by numerical simulations that highlight the strong dependence of spatial correlations on underlying network structure, we first analyze spatial correlations in infection dynamics on a series of regular networks with increasingly randomized connectivity. Focusing specifically on the limiting case of the regular random graph (RRG), we then generalize existing lattice-based theory, deriving a system of ODEs to model the temporal evolution of spatial correlations at varying shortest path lengths. Unlike traditional pairwise models—which only account for correlations between directly connected (nearest-neighbor) nodes—our approach resolves the full hierarchy of spatial correlations at every network distance by solving coupled ODEs for each shell around an infected node. This enables us to capture how infection clustering decays with distance and how finite network connectivity and the network’s structure shape epidemic trajectories, which are missed by standard closures. Comparison with numerical simulations demonstrates that our derived methodology yields substantially improved accuracy over mean-field predictions and lattice-based corrections. Furthermore, by systematically refining mean-field theory with spatial correction terms that account for local clustering, we derive closed-form asymptotic expansions for both the endemic infection density and the two-point correlation function in the long-time limit. These steady-state approximations not only quantify the impact of network structure on endemic persistence, but also provide expressions that show how finite connectivity suppresses infection prevalence relative to classical mean-field predictions. Thus, our results offer a robust analytical tool for understanding and forecasting the dynamics of infectious diseases in SIS processes on RRGs, revealing how structural randomness affects the spatial organization of infections and thereby shapes epidemic trajectories.
@@ II Corrections to SIS mean-field theory on regular lattices
$\frac{\mathrm{d}x_{I}}{\mathrm{d}t}=\gamma x_{I}-\beta k_{0}x_{I}x_{S},\;\;\;\frac{\mathrm{d}x_{S}}{\mathrm{d}t}=\beta k_{0}x_{I}x_{S}-\gamma x_{I},$ (1)
where $\beta>0$ is the rate of infection of one susceptible individual from one infectious individual, $\gamma>0$ is the recovery rate of each infectious individual, $k_{0}$ is the fixed degree of each node, and $0\leq x_{I},x_{S}\leq 1$ are the population fractions in the $I$ and $S$ states, respectively. Here, the ratio $R_{0}=\beta k_{0}/\gamma$ is known as the basic reproduction number, a summary statistic that, given a mean-field assumption, determines whether an infection will eventually vanish from a population or become endemic [13].
The conservation law $x_{S}+x_{I}=1$ means that every individual is either susceptible or infected at any point in time, and time is often rescaled as $t\to\gamma t$ by setting $\gamma=1$. Hence, Eq. (1) can be reduced to a single equation
@@
In Eq. (2) the first term represents the infection of $k_{0}$ susceptible neighbors by an infectious node, each at a rate $\beta$, while the second term represents the recovery of infectious individuals at a rate of $1$. Here, $x_{I}(1-x_{I})$ gives the probability that a randomly selected node is susceptible and its neighbor is infected, assuming, under the mean-field approximation, that all nodes share the same average infection probability, independent of their neighborhood. Yet, due to the localized nature of the infection, it is more likely that a node will be infected if its neighbor is infected as well, which significantly affects the predicted trajectory of a disease for low $k_{0}$, leading to inaccurate forecasting (red and black traces in Fig. 2).
The same limitations of mean-field theory appear in network structures with non-random connectivity, such as lattice networks. For this reason, many applications require refining mean-field approaches to achieve more reliable forecasting. Previous extensions to mean-field theory, especially in the context of birth-death-movement models, have established a strong foundation for accurately predicting population dynamics on lattices by systematically including spatial correlations into the analysis [11, 12]. Here, each lattice site can be either occupied $A$ or empty $\emptyset$. The allowed processes include: birth, where a cell can reproduce into an adjacent empty site $A+\emptyset\to A+A$; death, where a cell is removed, leaving an empty site $A\to\emptyset$; and movement, where a cell moves to a neighboring empty site $A+\emptyset\to\emptyset+A$. Of these, the birth and death processes can be directly mapped to infection $I+S\to I+I$ and recovery $I\to S$ events in SIS dynamics, while cell movement is omitted. In this context, corrections to mean-field theory for SIS processes can be derived under the assumption that the positions and connectivity of individuals remain invariant over time.
The infection process in network-based SIS models relies on transmissions occurring exclusively across edges connecting susceptible and infected individuals. This local interaction structure makes it advantageous to formulate the dynamics directly in terms of the density of susceptible-infected pairs, $x_{SI}^{(1)}$, rather than only global densities [8]. Here and below, the superscript $1$ denotes distance-one adjacent lattice sites. Thus, the evolution equation for the infected fraction can be written as
@@
where $x_{II}^{(r)}(t)$ and $x_{SI}^{(r)}(t)$ denote the fractions of infected-infected and susceptible-infected pairs at distance $r$, respectively. The parameter $r$ can refer to either Euclidean distance or shortest path length, but does not explicitly enter the system dynamics, as demonstrated in the following section. The normalization by mean-field edge densities, which assume statistical independence, ensures that $F_{II}^{(r)}$ and $F_{SI}^{(r)}$ quantify the degree of spatial correlation between node states at distance $r$ on the network.
Treating $x_{SI}$ and $x_{II}$ as edge probabilities, it follows that $x_{II}^{(r)}+x_{SI}^{(r)}=x_{I}^{(r)}=x_{I}$. This is since $x_{I}^{(r)}$ is the marginal probability of a node at distance $r$ from the reference node (regardless of its status) to be infected, and according to the mean-field assumption is equal to the fraction of infected individuals across the whole network. Therefore, the correlation functions are related as
@@
Using Eqs. (4) and (5) in Eq. (3), the pairwise model becomes an explicit correction to the mean-field equation,
$\frac{\mathrm{d}x_{I}}{\mathrm{d}t}=\beta k_{0}x_{I}\left[1-F_{II}^{(1)}\left(t\right)x_{I}\right]-x_{I}.$ (6)
This form illustrates how local clustering of infection, as captured by $F_{II}^{(1)}$, alters the epidemic trajectories predicted by mean-field theory.
The infected-infected pairwise correlation function $F_{II}^{(r)}\left(t\right)$ approaches unity when the infection states of nodes at distance $r$ are statistically independent, which occurs in well-mixed networks ($k_{0}\gg 1$), where infected nodes are expected to be randomly distributed without pronounced spatial clustering. Conversely, $F_{II}^{(r)}\left(t\right)$ is greater than unity when infected nodes exhibit clustering (positive spatial correlation), as commonly seen in network-based SIS processes due to local infection spread from neighbors, making infected nodes more likely to be found near each other than by chance. Under typical SIS dynamics on connected networks, values of $F_{II}^{(r)}\left(t\right)<1$ do not occur, since infection propagation inherently generates clustering rather than dispersion. Therefore, the term $1-F_{II}^{(1)}\!(t)x_{I}\;$ may be interpreted as an effective susceptible density, and is typically lower compared to the naive mean-field value. Similarly, the behavior of $F_{SI}^{(r)}(t)$ tends to lie below unity at intermediate distances due to local depletion of susceptible neighbors around infected sites. This attenuation reflects the saturation of infection fronts and limits the effective reach of contagion, beyond what the mean-field susceptible density would suggest.
Figure 2: Average infection density, $x_{I}$ (top row), and the distance-one pairwise correlation function, $F_{II}^{(1)}\left(t\right)$ (bottom row). Here, SIS model simulations on the RRG (red) are compared with numerical solutions of the corrections to mean-field theory derived in Section V (blue), and with mean-field theory (black). Simulations were executed on networks with $N=10^{4}$ nodes, with an initial infection density of $10^{-2}$ such that 100 infective individuals were placed randomly on the network, and with $\gamma=1$. The first two columns correspond to a node degree of $k_{0}=4$ with infection rates of $\beta k_{0}=1.7$ and $2.0$, while the last two columns correspond to $k_{0}=6$ with $\beta k_{0}=1.4$ and $1.7$, respectively. Results in each plot are averaged over $100$ simulations.

### v2

# Spatial correlations in SIS processes on random regular graphs

@@ Abstract
In network-based SIS models of infectious disease transmission, infection can only occur between directly connected individuals. This constraint naturally gives rise to spatial correlations between the states of neighboring nodes, as the infection status of connected individuals becomes interdependent. Although mean-field approximations and the standard pairwise model are commonly used to simplify disease forecasting on networks, they inadequately capture spatial correlations; mean-field frameworks assume that populations are well-mixed, while the pairwise model neglects correlations beyond nearest-neighbor connections, which leads to inaccurate predictions of infection numbers over time. As such, the development of approximations that account for higher order spatially correlated infections is of great interest, as they offer a compromise between accurate disease forecasting and analytic tractability. Here, we use existing corrections to mean-field theory on the regular lattice to construct a more general framework for equivalent corrections on random regular graph topologies. We derive and simulate a hierarchical system of ordinary differential equations for the time evolution of the spatial correlation function at various geodesic distances on random networks. Solving these equations allows us to predict the time-dependent global infection density, which agrees well with numerical simulations. Our results substantially improve on existing corrections to mean-field theory for infectious individuals in SIS processes and provide an in-depth characterization of how structural randomness in networks affects the dynamical trajectories of infectious diseases on networks.
@@ I Introduction
The Susceptible-Infected-Susceptible (SIS) model, rooted in the compartmental framework of Kermack and McKendrick [9], is a foundational framework for describing the spread of infectious diseases in which individuals do not acquire lasting immunity following infection (e.g. the common cold or many sexually transmitted diseases [2, 19, 3, 4]). In the model, members of a population transition repeatedly between two states: susceptible (S) and infected (I). The dynamics of the SIS model are governed by two processes: the infection per connection between infected and susceptible individuals, occurring at a rate $\beta$, and the recovery of infectious individuals at a rate $\gamma$ (see Fig. 1(a)). Many generalized SIS models assume that susceptible and infected populations are well-mixed spatially, such that the rates of contact between them are assumed to be uniform across a population. Such models are often referred to as mean-field approximations and, while they are highly amenable to mathematical analysis [18, 17, 3], they are generally inappropriate for real-world scenarios, where infected individuals are spatially ordered or clustered. In such contexts, the dynamical trajectory of a disease is highly sensitive to the spatial distribution of infected individuals, with the infection spreading more rapidly when infected individuals are distributed more evenly in space [7].
@@
Figure 1: (a) Schematic of the SIS model on networks. Infected nodes recover with rate $\gamma$ (conventionally, time is rescaled as $t\to\gamma t$, implying $\gamma=1$) and infect their nearest neighbors, each at rate $\beta$. (b)-(e) Snapshots of SIS model simulations on RRGs with $N=4000$ nodes in the steady state. The left column [(b), (d)] corresponds to a node degree $k_{0}=4$ and the right column [(c), (e)] to $k_{0}=50$; in addition, in (b) and (c) $\beta k_{0}=1.7$, and in (d) and (e) $\beta k_{0}=2$. Susceptible nodes are blue, whereas infectious nodes are colored by their local deviation from the mean-field expectation: red, gray, and green denote positive, zero, and negative spatial correlations, respectively. Stronger positive correlation is observed at lower $k_{0}$, while at higher $k_{0}$ a symmetric, uncorrelated mix of red and green fluctuations is apparent. Increasing the infection rate, $\beta$, also decreases these local correlations but also increases the overall infection density in the steady state. (f) Steady state infection density versus $k_{0}$, obtained from numerical simulations on RRGs with $N=10^{4}$ nodes, $\beta k_{0}=1.7$, and $\gamma=1$. The dashed line is the expected mean-field value of $1-1/(\beta k_{0})\simeq 0.41$. Each data point represents an average over 100 realizations. In all cases, 100100 infectious individuals are initially placed randomly on the network, with all other nodes being susceptibles.
One solution to mitigate the inaccuracies of mean-field models is to represent the SIS process on a network, in which nodes represent individuals in a population and edges represent connections between individuals (i.e., potential routes for the flow of infection). This explicit connectivity necessarily leads to the emergence of spatial correlations between infectious individuals, as transmission can only occur via direct links to currently infected neighbors (see example in Figs. 1(b)-(e)). This local transmission mechanism leads to epidemic trajectories that differ markedly from those predicted by mean-field models. This is especially evident on networks with lower average node degree, in which these deviations are more pronounced (Fig. 1(f)), or on regular lattices with a high degree of spatial order. A significant analytical advance in capturing these dynamics was the development of pairwise approximation methods. By tracking the densities of connected node pairs of various types (e.g., S-I pairs), these models successfully account for these emergent local correlations. Consequently, pairwise frameworks are more accurate than standard mean-field models, particularly in the aforementioned low-degree regimes and on heterogeneous topologies [7, 8, 3].
Despite these improvements, existing ODE-based approaches rely on closure approximations that assume spatial correlations between infectious individuals decay rapidly with shortest path distance [7, 1, 11]. The standard pairwise model (SPM) [7, 10] is inherently limited by this; it restricts its pair-level description exclusively to neighboring nodes, and neglects spatial correlations at larger distances between non-neighboring infected individuals. Thus, no analytical framework exists that systematically accounts for multi-scale spatial correlations on generic networks, and in particular, on random regular networks (RRGs), making analysis challenging and leading to a preference for simulation-based exploration.
@@
To address this challenge, this work extends the SPM to derive a general framework for the accurate forecasting of disease spread via SIS processes on RRG topologies. Motivated by numerical simulations that highlight the strong dependence of spatial correlations on underlying network structure, particularly at lower node degrees, we first analyze spatial correlations in infection dynamics on a series of regular networks with increasingly randomized connectivity. Focusing specifically on the limiting case of the RRG, we then generalize existing lattice-based theory, deriving a system of ODEs to model the temporal evolution of spatial correlations at varying shortest path lengths. Unlike the SPM, our multi-shell pairwise model (MPM) resolves the full hierarchy of spatial correlations at every network distance by solving a system of coupled ODEs describing the dynamics within each shell (i.e., the set of nodes at a given shortest-path distance) around an infected node. This enables us to capture how infection clustering decays with distance and how finite network connectivity and network structure shape epidemic trajectories, which are missed by standard closures. Comparison with numerical simulations demonstrates that our derived methodology yields substantially improved accuracy over mean-field and SPM predictions. Furthermore, by systematically refining mean-field theory with spatial correction terms that account for local clustering, we derive closed-form asymptotic expansions for both the endemic infection density and the pairwise spatial correlations in the long-time limit. These steady-state approximations not only quantify the impact of network structure on endemic persistence, but also provide expressions that show how finite connectivity suppresses infection prevalence, thereby reducing the accuracy of classical mean-field predictions. Thus, our results offer a robust analytical tool for understanding and forecasting the dynamics of infectious diseases driven by SIS processes on RRGs, revealing how structural randomness affects the spatial organization of infections and thereby shapes epidemic trajectories.
@@ II Corrections to SIS mean-field theory on regular lattices
$\frac{\mathrm{d}x_{S}}{\mathrm{d}t}=-\beta k_{0}x_{I}x_{S}+\gamma x_{I},\;\;\;\frac{\mathrm{d}x_{I}}{\mathrm{d}t}=\beta k_{0}x_{I}x_{S}-\gamma x_{I},$ (1)
where $\beta>0$ is the rate of infection of one susceptible individual from one infectious individual, $\gamma>0$ is the recovery rate of each infectious individual, $k_{0}$ is the fixed degree of each node, and $0\leq x_{I},x_{S}\leq 1$ are the population fractions in the $I$ and $S$ states, respectively.
The conservation law $x_{S}+x_{I}=1$ means that every individual is either susceptible or infected at any point in time, and time is often rescaled as $t\to\gamma t$ by setting $\gamma=1$. Hence, Eq. (1) can be reduced to
@@
In Eq. (2), the first term represents the infection at a rate $\beta$ of susceptible nodes by an average of $k_{0}x_{I}$ infected neighbors, while the second term represents the recovery of infectious individuals at a rate of $1$.
Under this mean-field assumption, $x_{I}(1-x_{I})$ gives the probability that an individual at a randomly selected node is susceptible and a specific neighbor is infected, which follows from the assumption that all nodes share the same average infection probability, independent of their location in the network neighborhood. However, due to the localized nature of the infection process, it is more likely that a node will be infected if its neighbor is infected as well. This significantly affects the predicted trajectory of a disease for low average node degree, $k_{0}$, leading to inaccurate forecasting using the mean-field approximation (dashed black traces in Fig. 2).
The same limitations of mean-field theory appear in network structures with non-random connectivity, such as lattice networks, where nodes are arranged in a highly ordered geometric grid and connect only to their immediate spatial neighbors. For this reason, many applications require refining mean-field approaches to achieve more reliable forecasting. Previous extensions to mean-field theory, especially in the context of birth-death-movement models, have established a strong foundation for accurately predicting population dynamics on lattices by systematically including spatial correlations into the analysis [1, 11]. In these works, each lattice site can be either occupied ($A$) or empty ($\emptyset$). The model permits three processes: birth, in which a cell reproduces into an adjacent empty site, $A+\emptyset\to A+A$; death, in which a cell is removed, leaving an empty site, $A\to\emptyset$; and movement, in which a cell moves to a neighboring empty site, $A\!+\!\emptyset\!\to\!\emptyset\!+\!A$. The birth and death processes map directly onto infection, $I+S\to I+I$, and recovery, $I\to S$, in SIS dynamics, while movement is neglected as we aim to derive corrections to mean-field theory for SIS processes.
The infection process in network-based SIS models relies on transmissions occurring only across edges connecting susceptible and infected individuals. This local interaction structure makes it advantageous to formulate the dynamics directly in terms of the density of connected susceptible-infected pairs, $x_{SI}^{(1)}$, rather than only global densities [7]. Henceforth, the superscript $1$ denotes (distance-one) adjacent lattice sites. Thus, the evolution of $x_{I}$ satisfies
@@
where $x_{II}^{(r)}(t)$ and $x_{SI}^{(r)}(t)$ denote the fractions of infected-infected and susceptible-infected pairs at a distance $r$ from one another, respectively. The parameter $r$ can refer either to Euclidean distance or shortest path length on the network, but does not explicitly enter the system dynamics, as demonstrated in the following section. The normalization by mean-field edge densities, which assume statistical independence (i.e., that the infection status of one node does not influence the probability of another being infected), ensures that $F_{II}^{(r)}$ and $F_{SI}^{(r)}$ quantify the degree of spatial correlation between node states at distance $r$ on the network.
Treating $x_{SI}$ and $x_{II}$ as edge probabilities, one has $x_{II}^{(r)}+x_{SI}^{(r)}=x_{I}^{(r)}=x_{I}$. This is since $x_{I}^{(r)}$ is the marginal probability of a node at distance $r$ from the reference node (regardless of its infection status) to be infected and, according to the mean-field assumption, is equal to the fraction of infected individuals across the whole network. Hence, the correlation functions are related as
@@
Plugging Eqs. (4) and (5) into Eq. (3), the pairwise model becomes an explicit correction to the mean-field equation,
$\frac{\mathrm{d}x_{I}}{\mathrm{d}t}=\beta k_{0}x_{I}\left[1-F_{II}^{(1)}x_{I}\right]-x_{I}.$ (6)
This equation, valid for homogeneous network topologies, illustrates how the local clustering of infectiousness, as captured by $F_{II}^{(1)}$, alters the prediction of the mean-field theory for the epidemic dynamics.
The infected–infected pairwise correlation function, $F_{II}^{(r)}$, approaches unity when the infection status of nodes at distance $r$ are statistically independent, which occurs in well-mixed networks ($k_{0}\gg 1$), where infected nodes are expected to be randomly distributed without pronounced spatial clustering. Conversely, $F_{II}^{(r)}$ is greater than unity when infected nodes exhibit clustering (positive spatial correlation), as commonly seen in network-based SIS processes due to local infection spread from neighbors, making infected nodes closer together than expected under independence. Under typical SIS dynamics on connected networks, values of $F_{II}^{(r)}<1$ do not occur, since infection propagation inherently generates clustering rather than dispersion. Therefore, the term $1-F_{II}^{(1)}\!x_{I}\;$ may be interpreted as an effective susceptible density, and it is typically lower compared to the naive mean-field approximation. Similarly, the behavior of $F_{SI}^{(r)}$ tends to lie below unity due to local depletion of susceptible neighbors around infected sites. This reduction reflects the rapid exhaustion of local susceptible pools, which suppresses the effective transmission rate beyond what the mean-field susceptible density would predict. Notably, although the SPM relies on an equation equivalent to Eq. (6) to describe infection dynamics, its formulation for the evolution of $F_{II}^{(1)}$ differs fundamentally (see Sec. IV). Consequently, as shown in Fig. 2, the SPM systematically overestimates the infection density at low average degrees $k_{0}$, whereas our MPM agrees well with corresponding numerical simulations.
Figure 2: Average infection density, $x_{I}$ (top row), and pairwise correlation function, $F_{II}^{(1)}$ (bottom row). Here, SIS model simulations on the RRG (red) are compared with numerical solutions of the MPM (blue), SPM (dark-green), and mean-field theory (dashed black). Simulations were executed on networks with $N=10^{4}$ nodes and $\gamma=1$, and in all simulations 100100 infectious individuals are initially placed randomly on the network, with all other nodes occupied by susceptible individuals, and all results are averaged over 100 simulations. The left column corresponds to $k_{0}=3$ with $\beta k_{0}=1.85$, while the right column corresponds to $k_{0}=4$ with $\beta k_{0}=1.7$.